Use the Rule of 72 to approximate when an investor can multiply his initial investment 75-fold, if the interest rate is 15 ½ %.
Answer #1

Rule of 72 says: for investment to become n times it takes =72*ln(n)/ln(2)*1/rate=72*LN(75)/LN(2)*1/15.5
=28.93386747 years